Math 291 Lab 4 Due Monday February 12, 2018

Instructions: Use L^{A}TEX to typeset a document containing each component described below. Turn in your lab by emailing it to fagerstrom@mnstate.edu. You should email both your raw TEX (.tex) file and your compiled document (in either .dvi or .ps or .pdf form).

You will be graded on both your raw TEX code and the accuracy of your compiled document. Name your file in the format of JaneDoeLab2.tex Include a name block similar to the one described in Lab 1

• Include the hyperref, multicol, and enumerate packages in your usepackage command.

• Then, start your built document with a link that links to a webpage of your choice (and one that I would be comfortable showing to my mother).

• Then typeset each of the following. Pay attention to the exact symbols used in each example. Some will require you to redefine the enumeration or itemize symbol used, and some will require multicolumns.

• If it is typeset on this sheet on its own page, then it should be typed on its own page in your submission as well. [The name block and hyperlink can share a page with the Billboard list.] A. Top 5 songs in 1976, 1986, 1996 and 2006 according to Billboard:

(a) 1976: i. Silly Love Songs by Wings ii. Don’t Go Breaking My Heart by Elton John and Kiki Dee iii. Disco Lady by Johnnie Taylor iv. December, 1963 (Oh What a Night) by The Four Seasons v. Play That Funky Music by Wild Cherry

(b) 1986:

i. That’s What Friends are For by Dionne Warwick, Gladys Knight, Elton John and Stevie

Wonder ii. Say You Say Me by Lionel Ritchie iii. I Miss You by Klymaxx iv. On My Own by Patti LaBelle and Michael McDonald v. Broken Wings by Mr. Mister

(c) 1996: i. Macarena (Bayside Boys Mix) by Los del Rio ii. One Sweet Day by Mariah Carey and Boyz II Men iii. Because You Loved Me by Celine Dion iv. Nobody Knows by the Tony Rich Project v. Always be My Baby by Mariah Carey

(d) 2006: i. Bad Day by Daniel Powter ii. Temperature by Sean Paul iii. Promiscuous by Nelly Furtado fet. Timbaland iv. You’re Beautiful by James Blunt v. Hips Don’t Lie by Shakira fet. Wyclef Jean I. I hope you dance II. Lose Yourself in Music.

a) Drift Away.

b) Turn the Radio Up. III. The heart of rock’n’roll is still beating. IV. Sing a Song

Goal of life #1: I hope you dance Goal of life #2: Lose Yourself in Music.

(aa) Drift Away.

(bb) Turn the Radio Up.

Goal of life #3: The heart of rock’n’roll is still beating. Goal of life #4: Sing a Song 1. This item has a space of 2.0 inches below it.

2. This item has a space of 1.5 inches below it.

3. This item has a space of 1.25 inches to the left of it.

4. This item has been moved 0.5 inches to the left and is writing on top of the counter.

(£) frog (£) lizard (£) chameleon (£) gecko (£) dragon (£) toad (£) salamander (£) newt (£) tadpole (£) crocodile (£) iguana (£) snake (£) pterodactyl

(£) eft